HUBBLE SPACE TELESCOPE DAILY REPORT #2279 PERIOD COVERED: 0000Z (UTC) 01/04/99 - 0000Z (UTC) 01/05/99 Daily Status Report as of 005/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Four Sets of WF/PC-2 7620 (WF/PC-2 Cycle 7 Standard Darks) The WF/PC-2 was used to make a series of Cycle 7 standard dark calibration images. This dark calibration program obtains dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed Eleven Sets of NIC/1/2/3 8093 (DARKs: Warming Up - Part II) The Near Infrared Camera (Cameras One, Two and Three) was used to make a series of dark calibration observations. The purpose of this proposal is to monitor the detector dark current during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed STIS/CCD/MA1 7481 (The Wind-momentum Luminosity Relationship for M31 and M33 B-supergiants) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to make observations of the B class supergiant star M-31-OB78-1 which is locate in the M-31 galaxy. The purpose of these observations is to acquire STIS low-resolution ultraviolet spectra of early B-supergiants in galaxies M31 and M33 to determine the terminal velocities of their stellar winds. The observations were completed as planned, and no anomalies were reported. 1.5 Completed Five Sets of NIC/1/2/3 8083 (FLATs: Warming Up - continuation) The Near Infrared Camera (Cameras One, Two and Three) was used to monitor the flat fields during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed Three Sets of WF/PC-2 7713 (WF/P-2 Cycle 7 Supplemental Darks pt3/3) The WF/PC-2 was used to make supplemental dark calibration images. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no anomalies were noted. 1.7 Completed Three Sets of WF/PC-2 7623 (WF/PC-2 Cycle 7 Internal Flats) The WF/PC-2 was used to make a set of internal Cycle 7 calibration flats which are designed to monitor the pixel to pixel flat field response and the visflat lamp degradation, as well as detect any possible changes due to contamination. The calibration observations were completed as planned, and no problems were reported. 1.8 Completed WF/PC-2 7377 (Dwarf Elliptical Galaxy Snapshot Survey II) The WF/PC-2 was used to make observations of the dwarf elliptical galaxy FCC246 as part of Snapshot Survey II. This survey will provide important information on the globular cluster systems of such galaxies and the properties of their cores and nuclei. The observations were completed as planned, and no anomalies were reported. 1.9 Completed Three Sets of STIS/MA1 7950 (Cycle 7 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA1) was used to make a dark calibration image. This is to establish if the detectors are performing nominally, that no internal movement of the MCP has taken place, and establish dark level in orbit for a given operating temperature. These darks will be used in later image reduction and also trended over instrument life time as indicators of detector degradation or damage. The observations were completed as planned, and no problems were reported. 1.10 Completed WF/PC-2 6699 (The Globular Cluster System of the Dwarf Elliptical System Galaxies NGC-185 and NGC 205) The WF/PC-2 was used to make observations of the globular cluster system around the nucleus of dwarf galaxy NGC-205. The propose is to carry out a general reconnaissance of the properties of the globular clusters. The observations were completed as planned, and no anomalies were reported. 1.11 Completed STIS/CCD 7912 (STIS Parallel Archive Proposal - Nearby Galaxies - Imaging and Spectroscopy) The Space Telescope Imaging Spectrograph (CCD) was used to make parallel observations of nearby galaxies. This survey will be useful to study the star formation histories, chemical evolution, and distances to these galaxies. These data will be placed immediately into the Hubble Data Archive. The observations were completed as scheduled, and no problems were reported. 1.12 Completed Eight Sets of STIS/CCD 8091 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation III) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude, non-scripted, parallel observations. The observations were completed as planned, and no anomalies were noted. 1.13 Completed Three Sets of STIS/CCD 8084 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make non-scripted parallel high galactic latitude observations as part of POMS test proposal. The observations were executed as scheduled, and no anomalies were noted. 1.14 Completed WF/PC-2 6355 (A UV Atlas of Nearby Galaxies) The WF/PC-2 was used to make ultraviolet observations of the nearby galaxy NGC-1507. The aim of the project is to build a reference ultraviolet atlas of galaxies whose optical images are well known, with the highest possible degree of information, covering all the morphological types and luminosity classes. The observations were executed as scheduled, and no anomalies were noted. 1.15 Completed NIC/2/3 7962 (NICMOS Warm-up Focus Monitoring) The Near Infrared Camera (Cameras Two and Three) was used to make observations of the globular cluster and nebula of NGC-3603 as part of the NICMOS warm-up focus monitoring. This proposal is used to determine and monitor the optimal focus settings for NIC/2 and NIC/3 during the dewar warm-up period at the end of 1998. The observations were executed as scheduled, and no anomalies were noted. 1.16 Completed WF/PC-2 6669 (Imaging of Two Dwarf Galaxies in Tidal Tails) The WF/PC-2 was used to make observations of tidal trails of the two merging dwarf galaxies in NGC-4038. These two dwarf galaxies are part of the M-81 group, and recent models suggest that it is interacting with, but has not formed from, the H-1 surrounding M81. The observations were completed as planned, and no anomalies were reported. 1.17 Completed Three Sets of NIC/2/3 8082 (NICMOS Pure Parallel Fall 98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel observations while another instrument is prime. The observations were executed as scheduled, and no anomalies were noted. 1.18 Completed NIC/2/3 8063 (NICMOS Pure Parallel Sponge - Backup) The Near Infrared Camera (Cameras Two and Three) was used to make observations taken in parallel with the prime science observations. This proposal is PRE-the new FOM default. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 5 Successful : 5 Scheduled Reacquisitions : 10 Successful : 10 2.2 FHST Updates: Scheduled : 15 Successful : 15 2.3 Operations Notes: Using ROP SR-1, the SSR EDAC error counter was cleared five times. The NICMOS detector warm-up has begun, indicating the end of cryogen. The NICMOS dewar temperatures are increasing at a rate of approximately 2 degrees K every 12 hours. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Today, at approximately 005/1455Z, the new CCS ground system will begain commanding the HST. Continuation of normal science observations and calibrations. /DMH